Nine years ago a mother was three times as old as her son in eight years from now the sum of their ages will be 78. How old are they today?



1
Expert's answer
2021-06-03T13:56:19-0400

Let a mother be "x" years old and her son be "y" years today.

Given nine years ago a mother was three times as old as her son 


"x-9=3(y-9)=>x=3y-18"

Eight years from now the sum of their ages will be 78


"x+8+y+8=78"

Substitute


"3y-18+y+16=78"

Solve for "y"


"4y=80"

"y=20"

Then


"x=3(20)-18=42"

a mother is "42" years old and her son is "20" years today.